OpenStudy (anonymous):

Which is the easiest way to prove that "If A is a unitary matrix and \( A^{-1} \) is also unitary matrix"

OpenStudy (jamesj):

If A is unitary matrix, then \[ A^{-1} = A^\dagger \] where \( A^\dagger \) is the conjugate transpose. As \[ (A^\dagger)^\dagger = A \] you should have your result.

OpenStudy (anonymous):

nice dagger!

OpenStudy (anonymous):

duh it was sooo easy, thanks James.

OpenStudy (anonymous):

now tell me how could we prove that |det (A)|=1

OpenStudy (anonymous):

\[(A^\dagger)^\dagger A^\dagger= AA^\dagger=I\]just testing

OpenStudy (jamesj):

Ah, that's straight forward too: \[ \det(A^\dagger) = (\det{A})^* \] where * means conjugate Now as \( AA^\dagger = I \) \[ 1 = \det I = \det(AA^\dagger) = \det A . (\det A)^* \] \[ \implies |\det A|^2 = 1 \implies |\det A| = 1 \]

myininaya (myininaya):

pretty!

myininaya (myininaya):

whats the * mean there james?

myininaya (myininaya):

oops

OpenStudy (jamesj):

complex conjugate. Hence ... \[ \det A^\dagger = \det (A^T)^* = (\det A^T)^* = (\det A)^* \]

myininaya (myininaya):

nvm i didn't read that line

myininaya (myininaya):

ok perfect

OpenStudy (anonymous):

hmmm i am sure it is right, but why does \[|\det A|^2 = 1 \implies |\det A| = 1\]?

OpenStudy (jamesj):

because |det A| has to be a non-negative number

OpenStudy (jamesj):

In other words, |det A|^2 = 1 implies |det A| = +1 or -1. But it can't equal -1.

OpenStudy (anonymous):

oh right.

OpenStudy (anonymous):

I still can't follow how \( \det A . (\det A)^* = |\det A| ^2 \) ?

OpenStudy (jamesj):

In genreal, for any complex number z, \[ zz^* = |z|^2 \]

OpenStudy (anonymous):

Aw, right! and your first result I think that is consistent to this : http://en.wikipedia.org/wiki/Determinant#Square_matrices_over_commutative_rings_and_abstract_properties ?

OpenStudy (anonymous):

That's a messy one though, do you know any easy proof of that property?

OpenStudy (jamesj):

which property in particular? Do you mean that \[ A^\dagger = A^{-1} \] or something else?

OpenStudy (anonymous):

That can be deduced easily from the definition of unitary matrices, I was wondering about \( \det(A^\dagger) = (\det{A})^* \) In words, the determinant of the complex conjugate of a complex matrix (which is also the determinant of its conjugate transpose) is the complex conjugate of its determinant, right ?

OpenStudy (jamesj):

oh, that's what I wrote out above. dagger = transpose and conjugate. Now the det of a transpose of a matrix is equal to the det of the original matrix. So det of A-dagger = the conjugate of det A.

OpenStudy (anonymous):

yes yes, but how to prove it ?

OpenStudy (jamesj):

Carefully, \[ \det[ A^\dagger ] = \det [ (A^T)* ] \ \ \ \hbox{by definition} \] \[ = (\det[ A^T ])^* \] because the conjugate of a sum of product of complex numbers is the sum of the product of the conjugate of each number, because (z+w)* = z* + w* and (zw)* = z*w* \[ = (\det A)^* \ \ \ \ \hbox{ because } \det A^T = \det A \ \hbox{ for all square matrices } \]

OpenStudy (anonymous):

Cool! should have reviewed complex numbers once, Thanks james.